# Agent Design Pattern Catalogue: A Collection of Architectural Patterns for Foundation Model based Agents

### v1

@@ 4.5 One-Shot Model Querying
Known uses: One-shot model querying can be considered configuration or use by default when a user is leveraging a foundation model, while CoT and Zero-shot-CoT both exemplify this pattern [25].
@@ 4.6 Incremental Model Querying
  • Reasoning uncertainty. The reasoning process may be oversimplified and hence endure uncertainties with only one attempt of model querying.
@@
  • HuggingGPT8. The underlying foundation model of HuggingGPT is queried multiple times to decompose users’ requests into fine-grained tasks, and then determine the dependencies and execution orders of tasks [26].
@@ 4.7 Single-Path Plan Generator
Context: An agent may not be proficient at performing particular tasks, while users also care about the process of how an agent produces the outputs.
@@
  • Reasoning uncertainty. Users may assign tasks with high-level abstraction, which may be challenging for agents to handle the uncertainty or ambiguity in the provided context.
@@
Solution: Fig. 9 illustrates a high-level graphical representation of single-path plan generator. After receiving and comprehending users’ goals, the single-path plan generator can coordinate the creation of intermediate steps for other agents or tools, to progressively lead towards goal accomplishment. Specifically, each step in this process is designed to have only a single subsequent step, such as Chain-of-Thought (CoT) [29]. Self-consistency is employed to confirm with the foundation model several times and select the most consistent answer as the final decision [30].
@@
  • Reasoning certainty. Single-path plan generator generates a multi-step plan, which clarifies the reasoning process and mitigate the uncertainty or ambiguity for achieving users’ goals.
@@ 4.8 Multi-Path Plan Generator
Context: An agent may not be proficient at performing particular tasks, while users also care about the process of how an agent produces the outputs.
@@
  • Reasoning uncertainty. Users may assign tasks with high-level abstraction, which may be challenging for agents to handle the uncertainty or ambiguity in the provided context.
@@ 4.9 Self-Reflection
  • Reasoning uncertainty. There may be inconsistencies or uncertainties embedded in the agent’s output, affecting the response accuracy.
@@
  • Continuous improvement. The agent can continuously update the knowledge base and the manner of formalising the prompts and knowledge, to provide more reliable and coherent output to users without or with fewer reflection steps.
@@ 4.12 Voting-based Cooperation
Solution: Fig. 12 illustrates how agents can cooperate to finalise a decision via votes. Specifically, an agent can first generate a candidate response to the user’s prompts. The agent queries a coordinator agent for reflective feedback, while the latter agent holds a vote in which different reflective suggestions are presented as choices. Additional agents are requested to submit their votes to select the most appropriate feedback according to their capabilities and experiences. The voting result will be formalised and sent back to the original agent, who can refine the response accordingly before answering the user.
@@
  • Accountability. The overall process and file results are recorded in the respective voting system. Stakeholders can trace back to identify the accountable agents selecting certain options.
  • Response accuracy. The finalised decisions after votes can leverage the strengths of multiple agents (e.g. comprehensive knowledge base), hence they are regarded as more accurate and reliable than the ones generated by a single agent.
@@ 4.13 Role-based Cooperation
Solution: Fig. 13 illustrates a high-level graphical representation of role-based cooperation. In particular, an agent-as-a-planner can generate a multi-step plan by decomposing user’s goal into a chain of tasks. Subsequently, the agent-as-an-assigner can orchestrate task assignment, i.e., some tasks can be completed by the assigner itself, while other tasks can be delegated to specific agent-as-a-worker based on their capabilities and expertise. In addition, if there is no available agent, agent-as-a-creator can be invoked to create a new agent with a specific role, by providing necessary resources, clear objectives and initial guidance to ensure a seamless transition of tasks and responsibilities.
@@
  • XAgent17. XAgent consists of three main parts: planner agent for task generation, dispatcher agent for task assignment, and tool agent for task completion.
@@ 4.14 Debate-based Cooperation
Solution: Fig. 14 depicts a high-level graphical representation of debate-based cooperation. A user can send queries to an agent, which will then share the questions with other agents. Given the shared question, each agent generates its own initial responses, and subsequently, a round of debate will start between the agents. Agents will propagate their initial response to each other for verification, while also providing instructions and potential planning directions to construct a more comprehensive response based on inclusive and collective outcomes. In addition, agents may utilise a shared memory in certain circumstances, or allow each other to access the respective memory facilitating the debate. This debate process can be iterative to enhance the performance of all participating agents. Debate-based cooperation can end according to a predefined number of debate rounds, or the agents will continue the procedure until a consensus answer is obtained.
@@
  • Response accuracy. During the cooperation, agents can provide accurate and up-to-date knowledge to each other, ensuring the reliability of eventual responses.
@@ 4.15 Multimodal Guardrails
Solution: Fig. 15 presents a simplified graphical representation of multimodal guardrails. Guardrails can be applied as an intermediate layer between the foundation model and all other components in a compound AI system. When users send prompts or other components (e.g. memory) transfer any message to the foundation mode, guardrails can first verify whether the information meets specific predefined requirements, only valid information will be delivered to the foundation model. For instance, personally identifiable information should be treated with care or removed to protect privacy. Equivalently, when the foundation model creates results, the guardrails need to ensure that the responses do not include biased or irrespective information to users, or fulfil the particular requirements of other system components. Please note that a set of guardrails can be implemented where each of them is responsible for specialised interactions, e.g., information retrieval from datastore, validation of users’ input, external API invocation, etc. Meanwhile, guardrails are capable of processing multimodal data such as text, audio, video to provide comprehensive monitoring and control.
@@ 5 Discussion
Evaluation of foundation model-based agents. Evaluations on agents and the underlying foundation models are significant to ensure they behave as intended. UK AI Safety Institute devised an evaluation framework for large language models [^23] https://ukgovernmentbeis.github.io/inspect_ai/. Whilst, the majority of pattern benefits and drawbacks are software quality attributes, which still require quantification for fine-grained metrics and rubrics. For instance, accountability can be further divided into three criteria of responsibility, auditability, and redressability, and each criterion has its own process, resource, and product metrics [61]. Proper quantification can promote the evaluation of agents and validate the effectiveness of applied patterns. Please note that the evaluation of foundation model-based agents can be categorised into system-level and component-level [62], while each granularity has respective testing methods and benchmarking.
@@ 6 Conclusion
Foundation model based agents are gaining increasing attention in various domains to intellectualise and automate the business process. However, practitioners are troubled by architectural challenges to design agents. Our previous work demonstrates a reference architecture to present an overview of agent design [3], while in this study, we scrutinise the forces, solutions, and trade-offs of 16 patterns. The pattern catalogue is provided as a holistic guidance for architects to better design and develop foundation model-based agents. In our future work, we will study how to apply the pattern catalogue with existing patterns to preserve the trustworthiness of agents, and further explore the architecture decisions that are related to foundation model-based agents.

### v2

@@ 4.5 One-Shot Model Querying
Known uses: One-shot model querying can be considered configuration or use by default when a user is leveraging a foundation model, while CoT and Zero-shot-CoT both exemplify this pattern [33, 34].
@@ 4.6 Incremental Model Querying
  • Oversimplification. The reasoning process may be oversimplified and hence endure uncertainties with only one attempt of model querying.
@@
  • HuggingGPT9. The underlying foundation model of HuggingGPT is queried multiple times to decompose users’ requests into fine-grained tasks, and then determine the dependencies and execution orders of tasks [35].
@@ 4.7 Single-Path Plan Generator
Context: A agent is considered “black box” to users, while users may care about the process of how an agent achieve users’ goals.
@@
  • Underspecification. Users may assign tasks with high-level abstraction, which may be challenging for agents to handle the uncertainty or ambiguity in the provided context.
@@
Solution: Fig. 9 illustrates a high-level graphical representation of single-path plan generator. After receiving and comprehending users’ goals, the single-path plan generator can coordinate the creation of intermediate steps for other agents or tools and prioritise the tasks, to progressively lead towards goal accomplishment. Specifically, each step in this process is designed to have only a single subsequent step, such as Chain-of-Thought (CoT) [38]. Self-consistency is employed to confirm with the foundation model several times and select the most consistent answer as the final decision [39]. Please note that the generated plan may have different granularity based on the given goal that complex plan may incorporate multiple workflows, processes, tasks and fine-grained steps.
@@
  • Reasoning certainty. Single-path plan generator generates a multi-step plan, which can reflect the reasoning process and mitigate the uncertainty or ambiguity for achieving users’ goals.
@@ 4.8 Multi-Path Plan Generator
Context: A agent is considered “black box” to users, while users may care about the process of how an agent achieve users’ goals.
@@
  • Underspecification. Users may assign tasks with high-level abstraction, which may be challenging for agents to handle the uncertainty or ambiguity in the provided context.
@@ 4.9 Self-Reflection
  • Reasoning uncertainty. There may be inconsistencies or uncertainties embedded in the agent’s reasoning process, affecting the task success rate and response accuracy.
@@
  • Continuous improvement. The agent can continuously update the memory or knowledge base and the manner of formalising the prompts and knowledge, to provide more reliable and coherent output to users without or with fewer reflection steps.
@@ 4.11 Human Reflection
  • Underspecification. Users may provide limited or ambiguous reflective feedback to agents.
@@ 4.12 Voting-based Cooperation
Solution: Fig. 12 illustrates how agents can cooperate to finalise a decision via votes. Specifically, an agent can first generate a candidate response to the user’s prompts, then it holds a vote in which different reflective suggestions are presented as choices. Additional agents are requested to submit their votes to select the most appropriate feedback according to their capabilities and experiences. In this circumstance, agents communicate in a centralised manner that the original agent will act as a coordinator. The voting result will be formalised and sent back to the original agent, who can refine the response accordingly before answering the user.
@@
  • Accountability. The overall procedure and final results are recorded in the respective voting system. Stakeholders can trace back to identify the accountable agents selecting certain options.
  • Collective intelligence. The finalised decisions after votes can leverage the strengths of multiple agents (e.g. comprehensive knowledge base), hence they are regarded as more accurate and reliable than the ones generated by a single agent.
@@ 4.13 Role-based Cooperation
Solution: Fig. 13 illustrates a high-level graphical representation of role-based cooperation, where agents coordinate in a hierarchical scheme. In particular, an agent-as-a-planner can generate a multi-step plan by decomposing user’s goal into a chain of tasks. Subsequently, the agent-as-an-assigner can orchestrate task assignment, i.e., some tasks can be completed by the assigner itself, while other tasks can be delegated to specific agent-as-a-worker based on their capabilities and expertise. In addition, if there is no available agent, agent-as-a-creator can be invoked to create a new agent with a specific role, by providing necessary resources, clear objectives and initial guidance to ensure a seamless transition of tasks and responsibilities. Please note that more elaborate roles can be defined and assigned to the agents.
@@
  • XAgent18. XAgent consists of three main parts: planner agent for task generation, dispatcher agent for task assignment, and tool agent for task completion.
@@ 4.14 Debate-based Cooperation
Solution: Fig. 14 depicts a high-level graphical representation of debate-based cooperation. A user can send queries to an agent, which will then share the questions with other agents. Given the shared question, each agent generates its own initial responses, and subsequently, a round of debate will start between the agents. Agents will propagate their initial response in a decentralised manner to each other for verification, while also providing instructions and potential planning directions to construct a more comprehensive response based on inclusive and collective outcomes. In addition, agents may utilise a shared memory in certain circumstances, or allow each other to access the respective memory facilitating the debate. This debate process can be iterative to enhance the performance of all participating agents. Debate-based cooperation can end according to a predefined number of debate rounds, or the agents will continue the procedure until a consensus answer is obtained.
@@
  • Data privacy. Agents may need to withhold certain sensitive information, which can affect the debate process.
@@ 4.15 Multimodal Guardrails
Solution: Fig. 15 presents a simplified graphical representation of multimodal guardrails. Guardrails can be applied as an intermediate layer between the foundation model and all other components in a compound AI system. When users send prompts or other components (e.g. memory) transfer any message to the foundation mode, guardrails can first verify whether the information meets specific predefined requirements, only valid information will be delivered to the foundation model. For instance, personally identifiable information should be treated with care or removed to protect privacy. Guardrails can evaluate the contents either relying on predefined examples, or in a “reference-free” manner. Equivalently, when the foundation model creates results, the guardrails need to ensure that the responses do not include biased or irrespective information to users, or fulfil the particular requirements of other system components. Please note that a set of guardrails can be implemented where each of them is responsible for specialised interactions, e.g., information retrieval from datastore, validation of users’ input, external API invocation, etc. Meanwhile, guardrails are capable of processing multimodal data such as text, audio, video to provide comprehensive monitoring and control.
@@ 4.16 Tool/Agent Registry
4.17 Agent Evaluator
Summary: Agent evaluator can perform testing to assess the agent regarding diverse requirements and metrics.
Context: Within an agent, the underlying foundation model and a series of components coordinate to conduct reasoning and generate the responses given users’ prompts.
Problem: How to assess the performance of agents to ensure they behave as intended?
Forces:
  • Functional suitability guarantee. Agent developers need to ensure that a deployed agent operates as intended, providing complete, correct, and appropriate services to users.
  • Adaptability improvement. Agent developers need to understand and analyse the usage of agents in specific scenarios, to perform suitable adaptations.
Figure 17: Agent evaluator.
Solution: Fig. 17 presents a simplified graphical representation of agent evaluator. Developers can deploy evaluator to assess the agent regarding responses and reasoning process at both design-time and runtime. Specifically, developers need to build up the evaluation pipeline, for instance, by defining specific scenario-based requirements, metrics and expected outputs from agents. Given particular context, the agent evaluator prepares context-specific test cases test cases (either searching from external resources or generating by itself), and performs evaluation on the agent components respectively. The evaluation results provide valuable feedback such as boundary cases, near-misses, etc., while developers can further fine-tune the agent or employ corresponding risk mitigation solutions, and also upgrade the evaluator based on the results.
Consequences:
Benefits:
  • Functional suitability. Agent developers can learn the agent’s behavior, and compare the actual responses with expected ones through the evaluation results.
  • Adaptability. Agent developers can analyse the evaluation results regarding scenario-based requirements, and decide whether the agent should adapt to new requirements or test cases.
  • Flexibility. Agent developers can define customised metrics and the expected outputs to test a specific aspect of the agent.
Drawbacks:
  • Metric quantification. It is difficult to design quantified rubrics for the assessment of software quality attributes.
  • Quality of evaluation. The evaluation quality is dependent on the prepared test cases.
Known uses:
  • Inspect [^22] https://ukgovernmentbeis.github.io/inspect_ai/. UK AI Safety Institute devised an evaluation framework for large language models that offers a series of built-in components, including prompt engineering, tool usage, etc.
  • DeepEval [^23] https://docs.confident-ai.com/. DeepEval incorporates 14 evaluation metrics, and supports agent development frameworks such as LlamaIndex, Hugging Face, etc.
  • Promptfoo [^24] https://github.com/promptfoo/promptfoo. Promptfoo can provide efficient evaluation services with caching, concurrency, and live reloading, and also enable automate scoring based on user-defined metrics.
  • Ragas [^25] https://github.com/explodinggradients/ragas. Ragas facilitates evaluation on the RAG pipelines via test dataset generation and leveraging LLM-assisted evaluation metrics.
Related patterns: Agent evaluator can be configured and deployed to assess the performance of other pattern-oriented agent components during both design-time and runtime.
@@ 5 Discussion
Evaluation of foundation model-based agents. Evaluations on agents and the underlying foundation models are significant to ensure they behave as intended. The majority of pattern benefits and drawbacks are software quality attributes, which still require quantification for fine-grained metrics and rubrics. For instance, accountability can be further divided into three criteria of responsibility, auditability, and redressability, and each criterion has its own process, resource, and product metrics [70]. Proper quantification can promote the evaluation of agents and validate the effectiveness of applied patterns. Please note that the evaluation of foundation model-based agents can be categorised into system-level and component-level [71], while each granularity has respective testing methods and benchmarking.
@@ 6 Conclusion
Foundation model based agents are gaining increasing attention in various domains to intellectualise and automate the business process. However, practitioners are troubled by architectural challenges to design agents. Our previous work demonstrates a reference architecture to present an overview of agent design [5], while in this study, we scrutinise the forces, solutions, and trade-offs of 17 patterns. The pattern catalogue is provided as a holistic guidance for architects to better design and develop foundation model-based agents. In our future work, we will study how to apply the pattern catalogue with existing patterns to preserve the trustworthiness of agents, and further explore the architecture decisions that are related to foundation model-based agents.